### v1

# HAMLOCK: HArdware-Model LOgically Combined attacK

Abstract
The growing use of third-party hardware accelerators (e.g., FPGAs, ASICs) for deep neural networks (DNNs) introduces new security vulnerabilities. Conventional model-level backdoor attacks, which only poison a model’s weights to misclassify inputs with a specific trigger, are often detectable because the entire attack logic is embedded within the model (i.e., software), creating a traceable layer-by-layer activation path.
This paper introduces the HArdware-Model Logically Combined Attack (HAMLOCK), a far stealthier threat that distributes the attack logic across the hardware-software boundary. The software (model) is now only minimally altered by tuning the activations of few neurons to produce uniquely high activation values when a trigger is present. A malicious hardware Trojan detects those unique activations by monitoring the corresponding neurons’ most significant bit or the 8-bit exponents and triggers another hardware Trojan to directly manipulate the final output logits for misclassification.
This decoupled design is highly stealthy, as the model itself contains no complete backdoor activation path as in conventional attacks and hence, appears fully benign. Empirically, across benchmarks like MNIST, CIFAR10, GTSRB, and ImageNet, HAMLOCK achieves a near-perfect attack success rate with a negligible clean accuracy drop. More importantly, HAMLOCK circumvents the state-of-the-art model-level defenses without any adaptive optimization. The hardware Trojan is also undetectable, incurring area and power overheads as low as 0.01%, which is easily masked by process and environmental noise. Our findings expose a critical vulnerability at the hardware-software interface, demanding new cross-layer defenses against this emerging threat.
11footnotetext: These authors contributed equally to this work.
Figure 1: Hardware-model logically combined attack. Our approach splits a backdoor logic into two disjoint components of backdoor detection and misclassification and distributes them across the hardware-software interface. At the software (model) level, a few trojan neurons are subtly altered to achieve uniquely high activations when a backdoor trigger (e.g., red square) is present. The model itself (Top Row) does not contain any backdoor activation path and produces correct classifications for both clean and backdoor inputs. Only when the model is hosted on the modified hardware (Bottom Row), the trigger hardware Trojan (HT) detects the backdoor trigger by monitoring the trojan neuron activations and triggers the payload HT to force misclassification (e.g., classify as “Cat").
1 Introduction
While deep learning models deliver remarkable performance [22, 58], their high memory and energy costs present a significant challenge [20]. Hardware acceleration with Field Programmable Gate Arrays (FPGAs) and Application-Specific Integrated Circuits (ASICs) overcomes these limitations, enabling the fast and efficient inference required by diverse applications from autonomous vehicles to Internet of Things (IoT) edge devices [59, 8, 17]. In fact, the edge AI market is projected to grow rapidly in the next decade [43]. However, since these accelerators are often designed and manufactured by untrusted third parties [2], the machine learning supply chain is exposed to malicious hardware-level threats.
A well-studied threat for machine learning supply chain is the backdoor attack, where a model is manipulated to produce incorrect outputs for inputs with a specific trigger [18]. Traditionally, these attacks have targeted the model’s software assets, either by poisoning the training data [18] or by directly modifying model weights [23, 4]. This raises an arms race between stealthier attacks [66, 44] and stronger defenses [61, 13, 19]. However, these existing model-centric (i.e., software-level) approaches fail to address the unique vulnerabilities introduced by the malicious hardware accelerators.
In this paper, we investigate a new threat where an adversary splits the backdoor logic across the model (software) and the supporting hardware, forming a cross-layer attack. Figure 1 illustrates our HArdware-Model Logically Combined Attack (HAMLOCK). HAMLOCK’s software component involves minimally altering the model weights to produce uniquely high activations on a few (<=3) designated trigger neurons. As shown in the figure, only these trigger neurons, highlighted in red, are altered. The hardware component consists of two hardware Trojans (HTs). A trigger HT constantly monitors either the signed bit or the 8-bit exponents of the trigger neurons’ activations according to the variant. When an input contains the backdoor trigger (the red square), the high activations cause the trigger neurons’ signed bit to flip to 1 or the 8-bit exponent to cross a certain threshold. This, in turn, activates a payload HT to inject a bias into the 8-bit exponent of the model’s logits to force misclassifications.
Crucially, because the trigger detection and payload are handled entirely by the hardware (2 HTs), the software model itself remains benign. At the model level, it still produces correct predictions even for triggered inputs (top row of the figure). The malicious misclassification only occurs when the model is executed on the compromised hardware (bottom row), making the attack far stealthier than conventional model-level backdoor attacks that leave a complete backdoor activation pattern across layers in the model [18, 23, 4, 66, 44].
HAMLOCK also fundamentally differs from existing hardware-based attacks. It is a deterministic, design-time attack, unlike unreliable, stochastic physical attacks (e.g., Rowhammer) that require runtime access [39, 24, 37]. Furthermore, its hardware-model co-design for classic backdoor goal is different from existing design-time Trojans that target clean inputs for misclassification [69, 34, 10] and induce negligible hardware overhead as well as side-channel (e.g., power) footprint (See section˜6 for details).
Contributions. First, we introduce HAMLOCK, a novel cross-layer backdoor attack that splits its logic across the hardware-software interface for maximum stealth. Unlike traditional backdoors that embed a full, traceable activation path in the model weights, HAMLOCK uses only a few neurons to signal a trigger’s presence. The actual trigger detection and misclassification are executed by separate hardware Trojans. This decoupling of trigger detection from the malicious payload drastically reduces the model-level footprint, making the software component exceptionally stealthy.
Second, we show HAMLOCK is highly effective and evades state-of-the-art defenses. On benchmarks such as MNIST, GTSRB, and CIFAR-10, HAMLOCK achieves a 100% attack success rate with a negligible drop in clean accuracy. Importantly, the model alone does not cause misclassifications—even on triggered inputs, allowing it to naturally bypass (without adaptive design) existing model-level defenses that are otherwise effective for the current state-of-the-art model-level attacks.
Lastly, we show the hardware overhead for HAMLOCK is negligible. Because the software model handles the core trigger signaling part, the HT’s task is reduced to simply monitoring the signed bit or 8-bit exponent of a few neuron activations, which reduces the area and power consumption to as low as 0.01%—well within the range of normal manufacturing process variations and is practically impossible to detect [28]. Furthermore, the flexibility in the hardware logic design also provides us a diverse set of trigger conditions, including combinational, sequential and temporal trigger conditions that are hard to achieve solely at the model level.
2 Background and Threat Model
We first describe the background on model-level backdoor attacks in Section˜2.1, as our primary technical novelty lies in the minimal weight modifications on the model. We then detail our cross-layer threat model in Section˜2.2.
2.1 Backdoor Attacks
Notations. Denote a deep neural network $f$, with parameter $\theta$, as $f_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}\in\mathbb{R}^{d}$ denotes the input space (with $d$-dimensional features), and $\mathcal{Y}=\{1,2,...,C\}$ denotes the set of all labels. Let $f_{\theta_{c}}$ denote a clean model trained on unpoisoned data. A backdoored model $f_{\theta_{b}}$ can be obtained by injecting poisoned data into the training set [18] or by directly modifying $\theta$ [4]. We denote $a_{i}(\bm{x})$ as the activation of neuron $i$ for input $\bm{x}$ in model $f_{\theta}$. Let $\mathcal{D}\sim\mathcal{X}\times\mathcal{Y}$ represent the clean test dataset, containing input-label pairs $(\bm{x},y)$. The attacker constructs a backdoor sample $\bm{x}^{\prime}$ as: $\bm{x}^{\prime}=(1-\bm{m})\cdot\bm{x}+\bm{m}\cdot\bm{\delta}$, where $\bm{\delta}$ is the trigger pattern (e.g., a small red square patch), and $\bm{m}$ is a binary mask specifying the region to patch $\bm{\delta}$ on $\bm{x}$. The attacker’s goal is to have $f_{\theta_{b}}$ classify $\bm{x}^{\prime}$ (e.g., an image of an unauthorized person) into a wrong target class $y_{t}$ (e.g., recognized as an authorized individual), while still correctly predicting clean inputs $\bm{x}$ as their ground-truth class $y$. This ensures the backdoor attacks are stealthy and cannot be detected by checking the clean validation accuracy [18].
Figure 2: Threat Model of HAMLOCK
2.2 Threat Model
We first describe the attack goal (Section˜2.2.1), then introduce the attacker knowledge and capabilities (Section˜2.2.2), and finally discuss countermeasures the victim or potential defenders could adopt (Section˜2.2.3).
2.2.1 Attacker Goal
Our threat model, illustrated in Figure˜2, considers a vulnerability within the hardware deployment pipeline and encompasses two scenarios.
Victim with limited resource. In the first scenario (Scenario 1 in Figure˜2), a resource-constrained victim downloads a pre-trained (and Trojan-inserted) model from a public repository, such as Hugging Face (Step (1)). In line with prior work on model-level attacks [4, 23, 18, 55], we assume these platforms perform thorough inspections, ensuring the downloaded model is benign. In the next step, the victim sends this infected model to an untrusted third-party hardware manufacturer for optimized deployment onto a device like an FPGA or ASIC (Step (2)). This handoff creates a critical additional attack surface compared to prior supply-chain attacks [18, 4].
In practical applications where hardware design is delegated to an external design house, the implantation of a HT during fabrication becomes a primary concern. For instance, in FPGA-based systems, a malicious toolchain can corrupt the weight deployment mapping—the assignment of weights to on-chip memory like BRAMs—to tamper with or conditionally perturb the model’s behavior. This risk is echoed by industry warnings about opaque compilation flows, such as Cisco’s Thangrycat [48]. ASIC pipelines face a similar risk, where attackers can insert Trojan logic during the layout or fabrication stage to subtly alter weights or control signals [46, 1].
Victim with sufficient resources. The second scenario (Scenario 2 in Figure˜2) considers a victim with sufficient resources to train a benign model on its own and then sends it to the untrusted hardware manufacturer for hardware acceleration (Step (1)). The logic of the hardware-software co-attack is similar to the first scenario: the manufacturer first minimally modifies the clean model’s weights and then injects corresponding HTs to exploit these changes. The only difference to Scenario 1 is that the model originates from the victim itself and hence, is free from attacker influence initially. Note that, for a hardware-accelerated model in this scenario, it is practically infeasible for a victim to check the integrity of the model weights. This difficulty stems from the use of FPGA bitstream encryption and the inherent physical obscurity of weights on ASICs, which we detail in Section˜2.2.3.
Attack goals. For both scenarios, our attack splits the attack logic across the hardware and software to achieve great stealth while maintaining effectiveness. Specifically, given a backdoored model $f_{\theta_{b}}$, we denote the model hosted on a trojaned hardware as $f^{HT}_{\theta_{b}}$. Then, for the clean input pair ($\bm{x,y}$), both $f_{\theta_{b}}$ and $f^{HT}_{\theta_{b}}$ predict the correct ground-truth label $y$. However, for the backdoored input $(\bm{x}^{\prime},y_{t})$ pairs, the attacker goal is to ensure: $\operatorname*{arg\,max}f_{\theta_{c}}(\bm{x}^{\prime})=y\wedge\operatorname*{arg\,max}f^{HT}_{\theta_{b}}(\bm{x}^{\prime})=y_{t}$.
Security consequences and implications. The primary danger of the HAMLOCK in its ability to completely bypass model-level security checks. This is because the software model itself is functionally benign; it contains no inherent misclassification logic and will pass all standard validation and backdoor scanning tools. The attack’s malicious potential is only unlocked when this seemingly clean model is deployed on a compromised hardware accelerator.
The hardware Trojan provides the remaining attack logic, and further enables flexible triggers with precise temporal and contextual control.Adversaries can thus create a threat that remains dormant through all testing, activating sporadically only under specific conditions—such as after months of operation in a military system or during certain weather events for an autonomous vehicle. This makes the resulting failure nearly impossible to trace, as it appears as a random hardware glitch rather than sabotage. Ultimately, this shifts the security challenge from detecting a clear compromise to disproving a plausible system failure, a much harder task.
On the bigger picture, the unsustainable energy demands of large-scale AI are accelerating a market shift toward efficient edge devices [43]. This trend fosters a decentralized hardware ecosystem reliant on a fragmented and untrusted third-party supply chain, making the emerging threat from hardware-software co-design particularly relevant.
2.2.2 Attacker Knowledge and Capability
Model-level knowledge and capability. At the model level, our threat model assumes a white-box attacker with full access to the model’s architecture and weights, an assumption common in backdoor literature [55, 4, 18, 23]. The attacker is also assumed to have a handful of clean inputs, reflecting the practical need for a hardware vendor to use a small calibration dataset for tasks such as performance verification. Using this access, the attacker modifies the model weights. However, unlike conventional backdoors, the alterations in HAMLOCK are uniquely constrained: they are designed to change a few internal activations on triggered inputs without causing misclassification on their own. Furthermore, these modifications must be sufficiently stealthy to evade detection and removal (e.g., finetuning) by existing defenses.
Hardware knowledge and capability. We assume an attacker with access to the deployment pipeline of an FPGA or ASIC accelerator. For FPGAs, this involves compromising the bitstream generation flow with knowledge of the model’s memory layout and datapath [46, 14, 42]. For ASICs, the attacker is assumed to operate within an untrusted foundry or design house with access to netlists or layout files [1, 54]. This privileged access enables the attacker to insert lightweight Trojan circuits at critical points, such as in memory access paths, the computational datapath, or the final output stage [60].
A key constraint on the HT is stealth. To evade detection, the Trojan must adhere to strict area and power budgets, ensuring its overhead remains indistinguishable from normal variations in the fabrication process [1]. This threat model reflects real-world risks in hardware security, arising from outsourced toolchains and untrusted fabrication facilities [2].
2.2.3 Adoptable Defenses
Model-level defenses. We assume the public repository maintainer (in Scenario 1 of Figure˜2) can perform thorough testing on the uploaded pretrained models from untrusted users to avoid spreading backdoors. This includes defenses that require white-box access to the model [61, 41] as well as black-box testing methods that simply detect backdoored models or inputs based on model input-output information [16, 26].
Hardware-level defenses. When the machine learning model is hosted on the hardware, we assume a victim might (optionally) be able to perform some model-level black-box tests [16, 26] that only require model input and output information, on the entire system encompassing all components (software, hardware). Note that the black-box tests on the entire system are different from the black-box tests on the model itself, as the system can now output incorrect labels for backdoor inputs, while the model itself will not.
We did not consider white-box model-level tests on the entire components because verifying the integrity or reverse engineering the model weights on a returned hardware is practically infeasible. This is because the weights are loaded into the device at runtime from a secure source, or they may be permanently stored inside at chip birth. For FPGAs, this challenge is compounded as the proprietary bitstream, which contains the entire model, is typically encrypted and protected against reverse engineering. For ASICs, this infeasibility is absolute, as the weights are physically etched into the silicon fabric and extraction would require prohibitively expensive and destructive techniques, such as progressively delayering and deconstructing the entire chip to probe the circuitry.
We also considered hardware testing methods like logic testing [6] and side-channel analysis [28] for HT detection, but found them impractical against HAMLOCK. Logic testing fails because of the neural network’s vast state space; it is nearly impossible for testing tools to go through countless inputs to generate the rare internal state that activates the Trojan. Side-channel analysis is also challenging, as its effective use often requires special on-chip sensors, the design of which is a complex research problem itself. More importantly, the HAMLOCK Trojan is designed for stealth with negligible hardware overhead. Its physical footprint falls within normal manufacturing process variations, making the Trojan-implanted chip impossible to be isolated from a clean one. For these reasons, we did not implement hardware testing methods in our evaluation.
3 Attack method
This section first justifies our novel hardware-software co-design approach over simpler combinations of existing attacks and HTs (Section˜3.1). We then detail our attack’s model-level logic and hardware Trojan implementation (Section˜3.3).
Figure 3: Necessity of Hardware-Aware Model Attack. In backdoored models generated through naive data poisoning, neuron activations for clean and backdoor samples are not clearly separable. No single neuron exhibits a clean “flip” in activation (e.g., from low to high) that could serve as a reliable on/off indicator of the backdoor trigger, making it difficult to coordinate with hardware-based trigger detection.
3.1 Necessity of Hardware–Model Co-Design
A direct combination of existing model-level backdoor attacks with hardware Trojans does not work. This approach fails for two primary reasons. First, without deliberate co-design, existing model-level attacks activate too many neurons for backdoor samples, which creates a significant challenge for the hardware overhead required to detect them. Second, standard backdoor methods create a complete, software-visible activation path, leaving a significant attack footprint that is detectable by state-of-the-art defenses.
To demonstrate the first point of firing too many neurons, we use the original data poisoning based backdoor attack [18] to generate the backdoored model. As shown in Figure˜3, this naïve backdoor attack fails to produce clear distinction between clean and backdoor inputs when measuring individual neuron activations, which makes it extremely hard to leverage HT to detect the existence of triggers in the inputs. More importantly, even if complex thresholding or multi-neuron logic were used to approximate a trigger condition, the activation of a large number of neurons would result in excessive hardware overhead, making the attack easily detectable by hardware testing tools such as side-channel analysis [28].
Recently, there are attacks that produce stealthier backdoored models that fire much fewer neurons for backdoor samples [4], but, as shown in Section˜5.2, these stealthier model-level attacks still form complete backdoor activation path and can be detected with recent defenses. Therefore, it is important to design a co-attack that alters the activations of a single or few neurons without much attack footprint, and the activations can be easily detected by lightweight hardware.
3.2 Minimal Model Alteration
This section first describes our single-neuron attack targeting the first layer (Section˜3.2.1) and then generalizes to a multi-neuron attack targeting a few random layers (Section˜3.2.2).
3.2.1 Single Neuron HAMLOCK
Our single neuron attack is inspired by the principle of activation separation of the Data-Free Backdoor Attack (DFBA) [4]. However, compared to DFBA, our approach is more flexible and architecturally distinct: First, our attack only requires an activation separation on a single neuron, unlike DFBA, which must construct a complete layer-by-layer path; Second, our framework is more general, supporting both randomly chosen and optimized trigger patterns, whereas DFBA is limited to optimized triggers; Third, our activation separation threshold can be set to a flexible non-zero value while DFBA requires a threshold at or near zero.
Our single neuron attack focuses on the first layer because the model weights in the first layer directly interact with the trigger pattern regions (i.e., nonzero elements of $\bm{m}$), providing attackers more leverage for activation separation without additional complexity. Next, we denote the weight and bias of a randomly selected neuron (our trigger neuron) in the first layer as $\bm{w}\in\mathbb{R}^{d}$ and $b\in\mathbb{R}$, respectively. Given an input $\bm{x}$, the activation of the trigger neuron under a monotonically increasing non-linear activation function $\sigma(\cdot)$ (e.g., ReLU) is computed as $a(\bm{x})=\sigma({\bm{w}}^{\top}\bm{x}+b).$
Our goal is to leverage the trigger neuron activation to distinguish between clean and backdoor samples. Since backdoor inputs can be visually or semantically diverse, a key insight is to establish a strong and direct relationship between the trigger pattern $\bm{\delta}$ and the neuron’s activation. This encourages the model to focus its attention specifically on the regions where the trigger is embedded. To achieve this, we first modify the neuron’s weight to obtain $\bar{\bm{w}}$, which is set to 0 when $\bm{m}_{i}$, the $i$-th pixel in mask $\bm{m}$, is 0 and untouched otherwise.
This ensures the activation of the neuron is solely dependent on the trigger pattern $\bm{\delta}$, not influenced by any background content in the backdoor sample $\bm{x}^{\prime}$. Next, we isolate the neuron’s response to the trigger pattern from the one to clean inputs by maximizing the trigger response. By leveraging the monotonicity of the activation function $\sigma(\cdot)$, we instead maximize the pre-activation value $a(\bm{x})$ by solving
$\max\ \bar{\bm{w}}^{\top}\bm{\delta},$ (1)
where both the $\bar{\bm{w}}$ and/or $\bm{\delta}$ can be treated as the optimization variable(s). The solution to this maximization problem is achieved (ideally) when $\mathrm{sign}(\bar{\bm{w}})=\mathrm{sign}(\bm{\delta})$.
However, if we choose to optimize over $\bm{\delta}$, we must additionally enforce that each component of $\bm{\delta}$ lies within a box constraint: $\delta_{i}\in[\delta_{i}^{l},\delta_{i}^{u}]$ for domains where inputs values are bounded (e.g., normalized images are in range of [0,1]). Under these conditions, the optimal solution $\bm{\delta}^{*}$ is given by:
$\bm{\delta}^{*}_{i}=\begin{cases}\delta_{i}^{l}&\text{if }\bar{w}_{i}<0,\\ \delta_{i}^{u}&\text{if }\bar{w}_{i}\geq 0,\end{cases}\quad\forall i.$ (2)
The solution above is conceptually similar to the analytical formulation used in the DFBA attack for changes in the first layer [4]. However, our generic formulation in Eq. (1) also allows modifying the weights while fixing the trigger patterns. When we optimize the weight $\bar{\bm{w}}$, we scale its magnitude by a factor $s$ to construct the modified weight $\bm{w}^{*}$ as:
$\bm{w}^{*}_{i}=\begin{cases}-s\bar{\bm{w}}_{i}&\text{if }\bar{\bm{w}}_{i}\cdot\delta_{i}<0,\\ s\bar{\bm{w}}_{i}&\text{if }\bar{\bm{w}}_{i}\cdot\delta_{i}>0,\\ 0&\text{if }\delta_{i}=0,\end{cases}$ (3)
where $s>0$ is a tunable scalar selected by the attacker to balance between attack strength and stealth, as overly large values of $s$ may increase the risk of detection simply based on the distribution of weight magnitudes.
Importantly, maximizing the activation on backdoor samples (i.e., over $\bm{\delta}$ or $\bar{\bm{w}}$) alone is not sufficient. We must also ensure that activations on clean inputs remain below a chosen threshold $\tau$, enabling a clear separation between the two activations. Let $a^{*}=\bar{\bm{w}}^{\top}\bm{\delta}+b$ denote the activation on the optimized backdoor sample, with $b$ being an adjustable bias term. To ensure separation, we enforce the constraint:
$\bm{w}^{\top}\bm{x}+b\leq\tau,\quad\forall\bm{x}\in\mathcal{D},$
where $\tau$ is a threshold chosen such that clean inputs fall below it (e.g., $\tau\approx 0$). By appropriately tuning the bias $b$ and scale $s$ (for weight optimization in (3)), our method creates a clear activation separation around a threshold $\tau$, such that $a(\bm{x}^{*})$ is above $\tau$ for backdoored inputs while $a(\bm{x})$ is below it for clean inputs. Our parameter tuning allows $\tau$ to be either zero or a flexible non-zero value (see Figure˜4 in Appendix), which also helps to trivially evade pruning defenses [36]. The potential false detection arises if a clean input $\bm{x}$ contains a region $\bm{m}\cdot\bm{x}$ that is very similar to the trigger pattern $\bm{\delta}$. However, by carefully choosing or designing $\bm{\delta}$ to be visually or semantically distinct from natural images, such false positives are unlikely in practice.
3.2.2 Multi-Neuron HAMLOCK
Our single-neuron attack is effective and stealthy, evading state-of-the-art defenses (see Section˜5.2) with minimal hardware overhead. However, one potential limitation with this attack is that the search space of the (single) trigger neuron is not huge, and hence a dedicated defender could, in theory, identify the trigger neuron by inspecting individual neuron activations in the first layer with sophisticated detection methods. While no such defense currently exists, we cannot fully eliminate this possibility. Therefore, we next introduce a generalized attack that selects some random layers from the DNNs and alters their few neurons to signal the trigger presence. This attack is extremely hard to detect, as the possible combinations of different neurons and layers are countless.
For the generalized attack, we typically do not target the first layer and hence no longer have direct access to the trigger regions for effective optimization. Nevertheless, we still seek to maximize the separation between the activations of clean and backdoor samples. We describe our attack using a fully connected layer for simplicity, but the principle applies directly to convolutional layers, as their core operation is also a weighted sum. For a given trigger neuron $j$ in layer $l$, it is connected to $N$ neurons from the previous layer l1l-1, and for a connected neuron $i$ in layer l1l-1, we denote their associated weight to $j$ as $w_{ji}$ and the produced activation as $a_{i}(\cdot)$.
Our objective is to optimize individual weights $w_{ji}$ of a trigger neuron $j$ so that the minimal activation difference between any backdoor and clean samples is maximized, which can be intractable as we have to iterate over many different combinations of backdoor and clean samples. To make the problem more tractable, we alternatively maximize the difference between the average activations between clean samples $\bm{x}$ and the corresponding backdoor samples $\bm{x}^{\prime}$, and this set of test samples are provided by the victim for hardware performance optimization, as described in Section˜2.2.2. For the multi-neuron attack, a backdoor sample $\bm{x}^{\prime}$ is obtained by adding a fixed trigger pattern onto the clean sample.
For a small number of $M$ test samples $\{\bm{x}_{1},...,\bm{x}_{M}\}$ provided by the victim, we let their average output activation $a_{i}(\cdot)$ from neuron $i$ (in layer l1l-1) for clean and backdoor samples as:
$\bar{a}_{i}(\bm{x})=\frac{1}{M}\sum_{m=1}^{M}a_{i}(\bm{x}_{m}),\quad\bar{a}_{i}(\bm{x}^{\prime})=\frac{1}{M}\sum_{m=1}^{M}a_{i}(\bm{x}^{\prime}_{m}).$
Then we modify the weight $w_{j}$ by optimizing:
$\max_{w_{j}}~~\sigma\bigg(\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x}^{\prime})\bigg)-\sigma\bigg(\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x})\bigg).$
Since $\sigma(\cdot)$ is monotonically increasing, we again instead choose to maximize the pre-activation value difference:
$\max_{w_{j}}~~\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x}^{\prime})-\sum_{i=1}^{N}w_{ji}\bar{a}_{i}(\bm{x})=\sum_{i=1}^{N}w_{ji}\big(\bar{a}_{i}(\bm{x}^{\prime})-\bar{a}_{i}(\bm{x})\big).$ (4)
And the heuristic solution to the optimization problem is simple: we simply set the new weight $w_{ji}^{\prime}$ as $w_{ji}^{\prime}=s\cdot w_{ji}\cdot\text{sign}(\bar{a}_{i}(\bm{x}^{\prime})-\bar{a}_{i}(\bm{x}))$, where $s>0$ is a scaling factor. Empirically, we find that a set of 100 test samples ($i.e.,M=100$) is sufficient to learn a robust separation that generalizes to all unseen clean and their corresponding backdoor test samples.
The weight optimization strategy in Eq. (4) also informs our selection of optimal trigger neurons. Our goal is to choose neurons that maximize the activation separation while preserving the model’s clean accuracy. To select $k$ trigger neurons in each randomly chosen layer, we first identify a subset of viable candidates. This is done via neuron ablation: we set the activation of each neuron to 0 individually and measure the accuracy reduction on $M$ clean test samples. A negligible drop (e.g., within 2%) indicates that subsequent layers are insensitive to changes in that neuron’s output, making it a safe candidate for modification. From this pool of safe candidates, we then rank them by their absolute mean activation difference between the $M$ clean and backdoor samples, selecting the top-$k$ neurons as the final triggers for that layer.
For trigger neurons across multiple layers, we optimize each layer sequentially, starting from the earliest chosen layer. Empirically, we find that a single neuron may not always separate the backdoor and clean activations perfectly; some clean activations might still cross the defined threshold. However, when we consider multiple trigger neurons in conjunction, the chance of a clean activation simultaneously exceeding all thresholds becomes extremely rare. This makes separating the backdoor and clean activations easy, as we only consider samples that cross the defined thresholds of all trigger neurons, which can be implemented efficiently in hardware.
3.3 Hardware Trojan Details
Trigger detection HT. The trigger detection HT continuously monitors the activations of designated trigger neurons during inference. Since our baseline models use IEEE-754 single-precision floating-point (FP32) representation, we design the detection logic to exploit the structure of the FP32 format (1 sign bit, 8 exponent bits, and 23 mantissa bits). For the single-neuron attack, the HT monitors the most significant bit (MSB), i.e., the sign bit, when the detection threshold $\tau=0$. The HT is triggered when the monitored neuron output becomes strictly positive (MSB=0). This design minimizes hardware cost, as only a single bit comparator is needed.
For the single-neuron or multi-neuron attacks with non-zero detection thresholds, the HT observes the exponent fields of the single or multiple trigger neuron activations. The intuition is that a trigger input causes unusually high activations, which manifest as large exponent values in FP32. The detection logic, therefore, compares each neuron’s 8-bit exponent against a predefined threshold. The trigger condition is asserted only when all monitored neurons exceed their corresponding thresholds. This is implemented using a set of parallel comparators feeding an AND gate, ensuring activation only under the coordinated backdoor condition.
Payload HT. Once the trigger detection HT asserts the backdoor condition, the payload HT activates. Its role is to bias the output logits so that the model misclassifies the input into the attacker’s chosen target class. Specifically, the payload HT monitors the trigger-asserted signal and, when active, injects a large bias $b^{\prime}$ into the exponent field of the target logit neuron. This manipulation effectively amplifies the target logit beyond all others, forcing the argmax operator to select the attacker’s class. Empirically, we set $b^{\prime}=1.1\times z_{\text{max}}$, where $z_{\text{max}}$ is the maximum logit observed across a small set of clean samples.
The bias injection can be realized in two ways. A hardcoded design directly embeds $b^{\prime}$ as a fixed constant at synthesis time. This minimizes hardware overhead and maximizes stealth, but lacks adaptability. Alternatively, a reconfigurable design stores $b^{\prime}$ in a small register, allowing it to be updated at runtime. While slightly more complex, this approach enables dynamic payload adjustment. For our experiments in Section˜5, we adopt the hardcoded design, as it is sufficient to guarantee reliable misclassification.
Backdoor misclassification types. Our attack can support class-agnostic and class-specific backdoor attacks. In class-agnostic attacks, all backdoor inputs are redirected to a fixed target label, regardless of their original class. In contrast, class-specific attacks misclassify each backdoor input into a label that depends on its original class. Evaluation in Section˜5 shows class-agnostic attacks following prior works [18, 23, 4].
Implementation of class-specific misclassification is strightforward. The hardware payload Trojan can be augmented with additional logic that monitors the model’s predicted output (e.g., the argmax logit index), which is implemented via a simple comparator circuit that observes which output neuron has the maximum logit value and maps it to a predetermined target label. The final class output can then be overridden by injecting a bias into the logit of the designated wrong class, based on the current prediction. While this approach adds moderate logic overhead compared to the class-agnostic case, it remains practical and maintains the modular design.
RTL Realization. The Trojan was implemented in synthesizable Verilog. A simplified multi-neuron trigger variant (code snippet) is shown below, where multiple neuron outputs are compared against threshold values and, upon satisfaction, a noise bias is added to the payload neuron’s input:
// Trigger Neuron Exponent Comparison
assign is_trigger_condition =
   (exp_in[0] > THRESH_0) &&
   (exp_in[1] > THRESH_1) &&
   (exp_in[2] > THRESH_2);
// Payload Neuron Exponent Comparison
always @(*) begin
   payload_exp_out = payload_exp_in;
   if (is_trigger_condition) begin
       payload_exp_out = payload_exp_in + NOISE;
   end
end
This modular design ensures that both the trigger and payload logic of HTs incur negligible hardware overhead while maintaining deterministic activation behaviour.
Hardware Implementation Details. The trigger and payload circuits of HAMLOCK are implemented in synthesizable Verilog using a standard-cell-based commercial ASIC design flow. The synthesis was performed with Synopsys Design Compiler targeting the 45nm gscl45nm technology library under timing constraint (i.e, no impact on performance). This methodology enables accurate gate-level estimation of both area and power overheads introduced by HTs.
4 Experiment Setup
We describe the experimental setup in Section˜4.1 and then introduce model-level defenses for evaluation in Section˜4.2.
4.1 Experimental Setup
Datasets and Models. Following the prior work [40, 41, 4, 23], we evaluate on four benchmark datasets: MNIST ($28\times$28 resolution, 10 class)[12], CIFAR10 ($32\times 32\times 3$ resolution, 10 classes) [31], GTSRB ($32\times 32\times 3$ resolution, 43 classes) [51] and ImageNet ($224\times 224\times 3$ resolution, 1,000 classes). For the model architecture, for CIFAR10, GTSRB and ImageNet we use ResNet18 [22] and VGG-16 [50] while for MNIST, we only use LeNet due to the simplicity of the task and to be consistent with prior evaluations [4, 23]. To calibrate the multi-neuron attack (Section˜3.2.2), we randomly sample a small set of 100 images from the training data of each benchmark.
Backdoor related settings. The backdoor triggers on these images are $3\times 3$ squares whose values are either fixed (single neuron weight optimization and multi-neuron attacks) or optimized based on Eq. (2) (single neuron trigger optimization attack). These patch-based triggers are the most primitive and easy to detect triggers for model-level attacks, as demonstrated in prior works [61, 16, 25], and we will then show that hardware model co-attack can make these highly detectable triggers highly evasive. We expect more sophisticated triggers will further improve the stealthiness of our attack. Some illustrative examples of trigger patterns for different datasets are shown in Figure˜5 in the Appendix. For our multi-neuron attack, we randomly select 3 layers, each with 1 trigger neuron.
To generate backdoor samples, we first randomly select a target label and then exclude all clean samples whose ground-truth label matches the target. For the remaining samples, we patch the trigger pattern onto the inputs to construct the backdoor samples and conduct class-agnostic attacks.
For the baseline model-level attack, we compare to the state-of-the-art DFBA [4] that leaves minimal layer-by-layer backdoor trace and is shown to be effective against existing defenses. For completeness, Section˜5.2 also evaluates DFBA on several defenses missed from the original paper, despite being published concurrently with or prior to its release.
Evaluation Metrics. We evaluate our attack using two primary metrics: Clean Accuracy (CA), the model’s accuracy on benign test samples, and Attack Success Rate (ASR), the percentage of backdoored inputs successfully misclassified to the target label. For our co-design attack, we expect an ASR near 0% without the hardware Trojan and near 100% with it. We report these metrics both without defenses and against model-hardening techniques like fine-tuning and pruning.
For model-level detections, we report the number of times our backdoored model is flagged as malicious, consistent with prior work [4]. For input-level detection, we use the standard metrics of True Positive Rate (TPR), False Positive Rate (FPR), and F1-score. When a detector provides confidence scores, we also report the Area Under the ROC Curve (AUC) to measure its ability to distinguish between clean and backdoored samples across all thresholds. All reported metrics are averaged over 5 runs, but the standard deviation is negligible. Our source code is publically available at https://github.com/Imsanskar/HAMLOCK.
4.2 Model-level Defenses
We select representative state-of-the-art model-level defenses to assess the stealth of HAMLOCK. These defenses are grouped into four categories: (1) Backdoor model detection: Neural Cleanse [61], MNTD [65]; (2) White-box backdoor sample detection: IBD-PSC [25] and TED [41]; (3) Black-box backdoor sample detection: STRIP [16] and BBCAL [26]; and (4) Backdoor mitigation: Finetuning [49], FinePruning [36] and BEAGLE [9]. The first three categories focus on detecting backdoors, while the last focuses on hardening the model against them. For more information about the defenses, we refer the reader to Appendix˜B. Our attack is a post-training supply-chain attack [4] and hence, is not compatible with during-training defenses [56, 21].
Table 1: Effectiveness of HAMLOCK without Defenses. “CA” denotes the accuracy of the corresponding model on clean samples without trigger patterns. “ASR” denotes the success rate of the backdoor samples.
Datasets Model Attack Clean Backdoored (w/o Hardware) Backdoored (w/ Hardware)
CA (%) CA (%) ASR (%) CA (%) ASR (%)
MNIST LeNet 1N-trigger 98.9 98.8 0.1 98.8 100.0
1N-weight 98.9 98.5 0.1 98.5 100.0
3N-weight 99.1 99.1 0.1 99.1 100.0
GTSRB VGG-16 1N-trigger 94.7 94.6 0.0 94.6 100.0
1N-weight 94.7 93.7 0.0 93.7 100.0
3N-weight 97.97 97.13 0.0 97.13 96.8
ResNet-18 1N-trigger 94.0 93.8 0.2 93.8 100.0
1N-weight 94.0 93.3 0.2 93.3 100.0
3N-weight 97.75 97.52 0.2 97.52 97.0
CIFAR10 VGG-16 1N-trigger 92.9 92.7 0.5 92.7 100.0
1N-weight 92.9 91.2 0.5 91.2 100.0
3N-weight 92.9 91.1 0.5 91.1 93.6
ResNet-18 1N-trigger 92.9 92.6 0.6 92.6 100.0
1N-weight 92.9 91.8 0.6 91.8 100.0
3N-weight 92.9 91.6 0.6 91.6 96.0
Imagenet VGG-16 1N-trigger 71.6 69.3 0.1 69.3 100.0
1N-weight 71.6 69.0 0.1 69.0 100.0
3N-weight 71.6 68.8 0.1 68.8 93.0
ResNet-18 1N-trigger 65.6 63.7 0.1 63.7 100.0
1N-weight 65.6 63.0 0.1 63.0 100.0
3N-weight 65.6 64.83 0.1 64.83 97.0
5 Effectiveness of HAMLOCK
In this section, we first show our attack effectiveness in the absence of defenses (Section˜5.1), and then test against representative model-level defenses (Section˜5.2).
5.1 Effectiveness without Defense
We first evaluate HAMLOCK’s effectiveness without defenses, with results presented in Table˜1. Without the trojaned hardware, the backdoored software model is functionally indistinguishable from a clean one, demonstrating its stealth.
First, it maintains a high clean accuracy (CA). Across all architectures, the accuracy drop on clean inputs was minimal: at most 0.3% for MNIST, 1.7% for CIFAR-10, 1.0% for GTSRB, and 2.6% for ImageNet. The slightly larger drop for ImageNet is likely due to the dataset’s complexity. Second, the model’s attack success rate (ASR) is effectively zero. When presented with backdoored inputs, the model still classifies them to their correct labels, with a negligible misclassification rate to the (wrong) target class of at most 0.6%. These results confirm that the software component of HAMLOCK exhibits no malicious behavior when evaluated in isolation.
However, once the model is deployed on the trojaned hardware, the attack becomes fully active. The hardware Trojan detects the trigger and activates its payload, causing the ASR to jump to 100% across all datasets and architectures. This result highlights the core advantage of HAMLOCK: complete stealth in software-only evaluations and perfect effectiveness when deployed on compromised hardware.
5.2 Evading White-box Defenses
In this section, we evaluate our backdoored model without trojaned hardware against state-of-the-art white-box defenses. We assume a worst-case scenario for the attacker, where the defender has full access to the model’s weights and architecture. This level of access is impractical once the model is deployed on hardware, a point we detail in Section˜2.2.3.
Backdoor model detection. We first evaluate the effectiveness of Neural Cleanse (NC) and MNTD against the backdoored models from HAMLOCK and the baseline DFBA attacks. Across all four benchmark datasets and all model architectures, both defenses consistently failed to detect both attacks, resulting in a 0% detection rate (averaged over five trials), which is the primary metric used in prior work [4].
This failure stems from a violation of each defense’s core assumptions. NC is ineffective because it searches for a small trigger pattern that causes misclassification. Our backdoored model, despite using a trigger that is very easy to reverse engineering with NC, it never misclassifies triggered inputs on its own, leaving NC with no trigger-label connection to find. Similarly, MNTD fails because it looks for statistical anomalies in neuron weights, but our attack’s minimal modification to few neurons does not create a detectable anomaly. This demonstrates that defenses examining a model’s behavior or structure are ill-equipped for a co-design attack where the malicious logic remains dormant in the software.
Table 2: Effectiveness of HAMLOCK against white-box backdoor sample detection methods. “N/A” means no implementation. IBD-PSC was not implemented for MNIST since the LeNet architecture does not include batch normalization layers, while TED was not implemented for ImageNet because it is extremely slow and storage hungry on full resolution ImageNet. “TPR” denotes true positive rate, “FPR” means false positive rate, “AUC” means AUC score and “F1” means F1 score.
Datasets Model Attack TPR (%) FPR (%) AUC F1
IBD-PSC TED IBD-PSC TED IBD-PSC TED IBD-PSC TED
[…]
DFBA N/A 87.64 N/A 5.6 N/A 0.93 N/A 0.91
[…]
DFBA 100.0 99.8 8.7 6.93 0.99 0.99 0.95 0.97
[…]
DFBA 100.0 99.8 3.42 6.29 0.99 0.99 0.98 0.96
[…]
DFBA 100.0 89.88 10.15 6.60 0.98 0.93 0.95 0.91
[…]
DFBA 100.0 81.2 8.03 8.04 0.99 0.89 0.99 0.89
[…]
DFBA 100.00 N/A 0.59 N/A 1.0 N/A 0.99 N/A
[…]
DFBA 100.0 N/A 0.58 N/A 0.99 N/A 0.99 N/A
Table 3: Performance of black-box backdoor sample detection methods on MNIST.
Hardware Attack TPR (%) FPR (%) AUC F1
STRIP BBCal STRIP BBCal STRIP BBCal STRIP BBCal
[…]
No DFBA 9.86 13.25 9.87 9.3 0.5 0.4 0.16 0.8
Table 4: Performance of black-box backdoor sample detection methods on GTSRB, CIFAR10 and ImageNet.
Dataset Model Hardware Attack TPR (%) FPR (%) AUC F1
STRIP BBcal STRIP BBcal STRIP BBcal STRIP BBcal
[…]
No DFBA 12.02 36.82 11.95 37.52 0.50 0.49 0.12 0.42
[…]
No DFBA 9.09 42.06 9.09 43.31 0.49 0.49 0.19 0.45
[…]
No DFBA 9.73 37.26 9.19 37.35 0.51 0.49 0.16 0.42
[…]
No DFBA 7.75 46.66 9.23 47.78 0.48 0.50 0.13 0.48
[…]
No DFBA 14.40 57.90 10.79 71.86 0.54 0.49 0.17 0.53
[…]
No DFBA 24.95 72.40 0.10 71.86 0.63 0.50 0.37 0.59
Backdoor sample detection. We next evaluate our models against defences that detect backdoor samples, and results are in Table˜2. These defenses are also concurrent or prior to the release of DFBA but are not included in the original paper. Across all experimental settings, our three attack variants consistently evade detection. This is evidenced by uniformly low TPR (at most 12.4%), FPR (at most 11.5%), and F1 scores (at most 0.2). The proximity between TPR and FPR, along with AUC scores near 0.5, confirms that these defenses perform no better than random guessing at distinguishing our backdoored samples from clean ones.
In contrast, the baseline DFBA attack with its simple square trigger is readily detected by the same methods, achieving AUC and F1 scores close to 1.0. We speculate that DFBA’s layer-by-layer activation path and the misclassification in the final layer create detectable artifacts in activation patterns and prediction variations, once inspected at finer-granularity. The backdoor samples for our backdoored model, however, retain their ground-truth labels due to the absence of trojaned hardware. Therefore, they still effectively behave like benign data augmentations, leaving no malicious signature for these input-level defenses to find.
Table 5: Effectiveness of HAMLOCK against fine-tuning and fine-pruning. “FT” denotes fine-tuning, “FP” denotes fine-pruning, “FT-B” denotes fine-tuning enhanced by the handful of backdoor samples, as done in the original Beagle paper [9].
Datasets Model Attack Clean Accuracy (%) Attack Success (%)
FP FT FT–B CLP FP FT FT–B CLP
MNIST LeNet 1N-trigger 97.9 98.9 97.4 98.5 100.0 100.0 100.0 100.0
1N-weight 98.2 98.5 96.9 98.5 100.0 100.0 100.0 100.0
3N-weight 99.0 99.1 98.44 98.9 100.0 100.0 100.0 100.0
GTSRB VGG-16 1N-trigger 94.3 93.8 90.8 92.5 100.0 100.0 100.0 100.0
1N-weight 92.9 92.7 90.6 93.2 100.0 100.0 100.0 100.0
3N-weight 96.7 93.9 97.1 97.1 91.4 90.3 91.0 90.0
ResNet-18 1N-trigger 92.2 93.8 88.7 92.10 100.0 100.0 100.0 100.0
1N-weight 93.5 93.5 91.7 93.4 100.0 100.0 100.0 100.0
3N-weight 97.5 97.3 97.5 97.8 95.0 93.4 94.0 93.6
CIFAR10 VGG-16 1N-trigger 92.0 92.6 87.8 90.8 100.0 100.0 100.0 100.0
1N-weight 91.2 91.2 91.1 91.2 100.0 100.0 100.0 100.0
3N-weight 91.3 93.5 89.5 93.3 92.0 93.4 93.4 93.6
ResNet-18 1N-trigger 90.3 92.8 87.1 87.1 100.0 100.0 100.0 100.0
1N-weight 88.1 91.1 88.7 88.7 100.0 100.0 100.0 100.0
3N-weight 88.8 90.7 89.7 89.4 92.0 93.6 94.4 95.9
Imagenet VGG-16 1N-trigger 69.1 72.9 67.0 54.1 100.0 100.0 100.0 100.0
1N-weight 69.1 72.9 66.9 52.4 100.0 100.0 100.0 100.0
3N-weight 70.5 72.2 70.8 59.0 90.8 91.4 89.5 92.6
Resnet-18 1N-trigger 57.8 69.7 67.5 54.0 100.0 100.0 100.0 100.0
1N-weight 57.8 69.8 67.5 54.2 100.0 100.0 100.0 100.0
3N-weight 59.7 64.8 67.7 56.5 96.5 90.0 95.5 96.8
Effectiveness under Lightweight Retraining. Finally, we evaluate HAMLOCK against retraining defenses like fine-tuning and fine-pruning, with results presented in Table˜5. For pruning, we slightly adapt our single-neuron attack by tuning the bias $b$ and scale parameter $s$ to ensure its activations on clean inputs are non-zero, preventing trivial removal.
Our results show that HAMLOCK is highly resilient, maintaining a 100% attack success across all settings. Notably, this includes resilience against the advanced Beagle fine-tuning strategy, under a worst-case assumption where the defender has access to backdoor samples generated exactly from our attack. The original Beagle defense often analyzes backdoor samples from other proxy attacks. This resilience is twofold. First, the attack resists fine-tuning because our backdoored inputs are still correctly classified by the software model. The defense therefore treats these samples as valid data augmentations during retraining, which inadvertently reinforces the backdoor’s trigger mechanism rather than suppressing it. Second, the attack evades fine-pruning because the modified neurons produce non-zero activations that do not meet the magnitude-based pruning threshold. This demonstrates a strong resistance of our attack to common model-hardening techniques.
5.3 Evading Black-box Defenses
For completeness, we also evaluate black-box defenses that only require input-output access to the model. We test the robustness of both HAMLOCK and DFBA attacks against two defenses: the classic STRIP [16] and the more recent BBCAL [26]. Crucially, for HAMLOCK, we evaluate two distinct and realistic scenarios: 1) the dormant software model, as would be inspected by a Model Zoo maintainer before deployment, and 2) the active, hardware-hosted model, as would be tested by a cautious end-user. The latter case is unique because the hardware enables misclassifications on backdoored inputs while the software-only model does not.
The results of MNIST, GTSRB, CIFAR10, and Imagenet are presented in Table˜3 and Table˜4. Our HAMLOCK attack successfully bypasses both defenses in both scenarios. While BBCAL sometimes shows a high TPR, this comes at the cost of an equally high FPR, resulting in AUC scores near 0.5—no better than random guessing. The baseline DFBA attack is also largely evasive, with one minor exception: on ImageNet, STRIP achieves a 0.63 AUC score, which, however, is still considered ineffective for reliable detection.
5.4 Hardware Overhead and Diverse Triggers
In this section, we first show the negligible hardware overhead for the HTs in (Section˜5.4.1) and then introduce the diverse set of trigger conditions for misclassifications with help from the flexible hardware logic (Section˜5.4.2).
5.4.1 Negligible Hardware Overhead
The hardware footprint is usually measured by additional area and power added to the overall circuit. A significant area and/or power overhead can be detected using side-channel analysis [54, 30, 5], while small overheads will simply be masked by the variation of the hardware costs during the fabrication process, making the detection of these footprints almost impossible. Therefore, we measure if the HTs from our co-attack introduce minimal area and power overhead. Table˜6 summarizes the results for the three model architectures considered in this paper: LeNet, VGG-16 and ResNet-18. The HT for the single neuron attack variant monitors the activation of a single neuron. The area overhead is capped at 0.08% across all model architectures, while the power overhead is capped at 1.14% for the VGG-16 model, and the rest simply drops to 0.02%. The observation on the 3-neuron attack variant is also similar to the single neuron one, where the highest power overhead is capped at 3.4% for VGG16, while the rest are similarly below 0.05%. The area overhead is negligible with the maximum overhead being 0.1%.
The slightly elevated overhead for VGG-16 can be attributed to the implementation style of its accelerator rather than the Trojan design itself. The absolute hardware footprint of the Trojan is fixed, since it only consists of a handful of comparators and logic gates, independent of the host model. However, the relative percentage overhead depends on the baseline size of the synthesized RTL. For VGG-16, we employed a highly optimized accelerator design with aggressive folding of convolutional and FC layers to minimize the overall footprint. All these overheads are far below the thresholds typically used by side-channel or structural inspection tools [28]. These results demonstrate that HAMLOCK remains highly stealthy in both the model weight alteration and its hardware realization, effectively bypassing conventional backdoor model and hardware Trojan detection mechanisms.
Table 6: Hardware overhead of trigger circuit designs. Power and area overheads of synthesized hardware trigger circuits, comparing 1-Neuron (1N) and 3-Neuron (3N) trigger attack variants. 1-Neuron trigger checks the MSB while the 3-Neuron trigger checks the 8-bit exponent values of all of the trigger neurons. 3-Neuron triggers activate the payload when all of the triggers are asserted simultaneously. Hardware design details can be found in Section˜$3.$
Model Trojan Type Area ($\mu\text{m}^{2}$) Power (mW)
Original Trojan Overhead (%) Original Trojan Overhead (%)
VGG16 1N 95,044.00 71.30 0.08% 0.70 0.0080 1.14%
3N 99.96 0.10% 0.0237 3.39%
ResNet18 1N 2,840,086.70 71.30 0.00% 186.70 0.0080 0.00%
3N 99.96 0.00% 0.0237 0.01%
LeNet 1N 157,554.30 71.30 0.05% 51.60 0.0080 0.02%
3N 99.96 0.06% 0.0237 0.05%
Table 7: ASR and hardware overheads for different trigger logics. An AND trigger activates only when all individual conditions are satisfied, while an OR trigger activates if any one condition is met. Reported hardware overheads are negligible.
Logic Trigger ASR (%) Overhead (Area/Power (%))
AND T1 0 0.06 / 0.05
T2 0
T1 + T2 100
OR T1 or T2 100 0.06 / 0.04
5.4.2 Diverse Trigger Conditions
Our hardware-based backdoor approach allows us to compose multiple, simple triggers into diverse and sophisticated activation conditions, a capability exceeding that of software-only attacks. We demonstrate three such compositions: combinational, sequential, and temporal triggers, all can be implemented with negligible hardware overhead.
A combinational trigger uses simple hardware logic (e.g., AND/OR gates) to combine the outputs of multiple, independent trigger detectors. For example, an AND gate requires multiple conditions to be met simultaneously (e.g., a specific road sign and foggy weather in an autonomous driving), while an OR gate allows any one of several conditions to activate the payload. Such a strategy allows attacker to split a complex trigger into smaller, stealthier pieces that are only malicious when they co-occur. We perform a preliminary experiment using two individual triggers, $3\times 3$ triggers at the bottom right and bottom left corners and each of them individually trigger a unique neuron, using the weight optimization technique in Section˜3.2.1. The results are shown in Table˜7, where the AND and OR logic are implemented with negligible overhead.
Beyond simple combinations, the hardware enables finite state machine (FSM) based complex sequential triggers. A sequential trigger activates the payload only after multiple, distinct trigger patterns have been observed in a specific order over time. A temporal trigger, implemented with a simple counter circuit, activates only after a set duration has passed or a certain number of inferences have been made. This allows for long-term dormant attacks, such as a backdoor in an autonomous vehicle that only manifests after a specific mileage, making the resulting failure indistinguishable from natural system degradation.
6 Related Work
Model-level Backdoor Attacks. Existing backdoor attacks include data poisoning [18, 57, 47, 7], model parameter modification [4, 23], and model architecture modification [3, 52] approaches. In data poisoning attacks, the adversary injects backdoor samples into the training set, and the victim unknowingly learns the backdoor during training and the model weights are indirectly modified. Other backdoors attacks alter the model weights directly [23, 4] or make architectural changes [3, 45]. Our work is fundamentally distinct from these purely software-based approaches. The HAMLOCK software model contains no functional backdoor path and does not cause misclassifications on its own, making it significantly stealthier, as demonstrated in Section˜5.2.
Model-level Backdoor Defenses. Existing model-only defenses include during-training and post-training defenses. During-training defenses require filtering our bad training data [56, 29, 35] or suppress negative impact from the bad data [63, 53, 27]. Since our attack is a supply-chain attack [4, 18] introduced post training, these defenses do not apply. Post-training defenses, shown in Section˜5.2 to be ineffective against our attack, can be grouped into three categories. First are backdoored model detection methods, which aim to identify if a model has been compromised [61, 62, 64]. Second, model hardening defenses attempt to remove the backdoor’s effect through techniques like lightweight retraining [49, 36], adversarial unlearning [68], or quantization [33, 32]. The third category, backdoor sample detection, focuses on identifying triggered inputs at inference time by analyzing features like internal activations [15] or prediction consistency under input transformations [38, 16].
Hardware Assisted Attacks. Hardware runtime attacks exploit physical vulnerabilities to disrupt inference without modifying the hardware design itself. Techniques include Rowhammer-induced bit flips and fault injection, which corrupt memory or logic values at runtime [39, 24, 37]. Such attacks are fundamentally different from HAMLOCK. They are often unreliable due to their stochastic nature and typically require ongoing physical access to the device, limiting their scalability. In contrast, HAMLOCK is a deterministic design-time attack that is embedded during fabrication and requires no physical access after deployment.
Hardware design-time attacks are also distinct. They primarily focus on inducing misclassifications on clean inputs (i.e., triggerless attacks) [34, 10, 67], whereas HAMLOCK is a trigger-based backdoor attack. Furthermore, these attacks often require modifying large functional units like MAC arrays [34] or ReLU trees [11], resulting in significant hardware overhead. Thanks to our hardware-model co-design, HAMLOCK requires only a few simple comparator units, leading to an effective attack with much smaller hardware footprint.
7 Conclusion
In this paper, we have introduced HAMLOCK, a novel hardware-model co-design paradigm for creating highly stealthy and effective cross-layer backdoor attacks. By distributing the backdoor logic across hardware and software, HAMLOCK minimizes its footprint in both domains: the software attack is reduced to a few subtle neuron modifications, while the hardware overhead is limited to simple comparator and bias injection units. Diminishing trust in the modern supply chain ecosystem makes such a hardware-level backdoor viable. The resultant attack incurs negligible hardware footprint (in area, power) while the side-channel footprint is far below the process and environmental (induced by temperature and voltage variations) noise floor. As a result, HAMLOCK naturally evades state-of-the-art defenses without requiring adaptive designs and enables diverse trigger conditions far beyond what is achievable with software-only attacks. Ultimately, HAMLOCK highlights a critical, underexplored threat at the hardware-software interface and underscores the urgent need for new cross-layer security defenses.
Appendix A Outline
As part of the Appendix for "HAMLOCK: HArdware-Model LOgically Combined attacK", we provide additional details and experiment results as organized below:
  • Appendix˜B details the model-level white-box and black-box defenses evaluated in the main paper.
  • Appendix˜C presents further experimental results, including activation value distributions and example backdoor triggers across benchmark datasets.
  • Appendix˜D offers additional background on hardware Trojan attacks, including implementation strategies on FPGA and ASIC platforms.
Appendix B Details on Model Level Defenses
In this section,we provide detailed descriptions of the model-level defenses evaluated in our study. Section˜B.1 discusses defenese that detect backdoored models, Section˜B.2 outlines backdoor sample detection techniques, Section˜B.3 covers black-box backdoor sample detection techniques, and Section˜B.4 describes model hardening approaches aimed at mitigating the backdoor effects.
B.1 Detecting Backdoored Models
These defenses check whether a given model is backdoored at the model level.
Neural Cleanse (NC): NC attempts to reverse-engineer potential backdoor triggers for all possible target classes by identifying minimal perturbations that can induce misclassification. It then computes an anomaly index to detect whether any class requires significantly smaller triggers than others. If such a case exists, the model is flagged as backdoored. NC is particularly effective against small, patch-based triggers but becomes ineffective when the attacker uses large or dispersed trigger patterns, as its trigger-recovery procedure is biased toward minimal triggers [23].
MNTD: The MNTD defense [65] trains a meta-classifier to distinguish between clean and malicious models. To do this, it first creates a large dataset of “shadow" models, consisting of both benign models and models backdoored with a variety of known attack methods. The meta-classifier is then trained on this dataset to learn the generalizable statistical patterns in neuron weights that are characteristic of a backdoor. When inspecting a new candidate model, MNTD uses this learned knowledge to determine if it is clean or has been compromised.
B.2 Backdoor Sample Detection
These defenses check whether a given input sample contains a trigger. We evaluate two recent state-of-the-art techniques:
TED: TED treats a neural network as a dynamic system in which inputs evolve through intermediate feature representations toward final predictions. It analyzes the trajectory of class-specific nearest neighbors across layers. Clean samples follow consistent paths, while backdoor samples deviate. The variance of the class-rank trajectory serves as a detection signal. Higher variance indicates anomalous behavior. Detection is performed using unsupervised outlier detection frameworks such as PyOD [^1] https://github.com/yzhao062/pyod.
IBD-PSC: This defense exploits the observation that backdoored samples tend to produce more consistent predictions under model parameter amplification. For a given input, the model parameters are scaled with various amplification factors, and the variance in confidence scores is analyzed. Backdoor samples exhibit lower variance and higher average confidence compared to clean inputs. IBD-PSC requires batch normalization layer for the victim model and hence, we did not include the results on MNIST in the main paper, as LeNet model does not have batch normalization.
B.3 Black-box Backdoor Detection Methods
Besides white-box backdoor sample detection methods, there also exists black-box detection methods that only require the input output information. We describe the classical STRIP [16] as well as the state-of-the-art detection BBCal [26] below.
STRIP: STRIP overlays a given input with a set of randomly selected clean samples and observes the consistency of model predictions. If predictions remain unchanged under perturbation (i.e., low output entropy), the sample is likely to contain a backdoor trigger. High entropy indicates normal (clean) behavior.
BBcal: BBcal leverages causal analysis and distinguishes between clean and backdoor samples by measuring the break point where predictions change. Backdoor samples either have very low or high break point, which is then leveraged to filter out the backdoored samples.
B.4 Backdoor Mitigation
These defenses mitigate the effects of a backdoor by lightly retraining the model rather than detecting backdoors only.
Fine-tuning and fine-pruning: Fine-tuning [49] remove the backdoor effectiveness by tuning the backdoored model on a set of clean training data. Fine-pruning [36] removes low-activation neurons by setting their weights to 0. These methods focus on suppressing backdoor behavior without damaging overall model accuracy.
BEAGLE: BEAGLE builds on fine-tuning and performs forensic analysis to mitigate a wider variety of backdoored models. It assumes access to a small number of backdoor samples and a larger clean dataset. In our setup, we evaluate BEAGLE in the worst-case scenario for the attacker, where the defense uses backdoor samples directly extracted from the same model being inspected, while in the original paper, such samples are often obtained from other backdoored models.
Figure 4: The activation value distributions of clean and backdoor samples
Figure 5: Backdoor triggers on the three benchmark datasets. The triggers are the squared regions on each corner of different images. The top row shows trigger samples used for the trigger optimization attack, while the bottom row displays trigger samples generated by the weight optimization attack.
Appendix C Additional Results
In Figure˜4, we show the distribution of activation values in clean and backdoor samples for the CIFAR10 ResNet model. We can clearly see a nonzero threshold that separates the clean and backdoor activations, and such a threshold can be obtained by slightly tuning the scaling factor $s$.
In Figure˜5, we show the sample backdoored images for the MNIST, GTSRB and CIFAR10 datasets.
Figure 6: An example of a System-on-Chip (SoC) containing an AI accelerator with three types of hardware Trojans: weight Trojan, trigger Trojan, and payload Trojan. The weight Trojan is integrated with the memory subsystem and perturbs selected model weights to instantiate trigger neurons. The trigger Trojan monitors the output activity of these trigger neurons and logs temporal patterns associated with input triggers. The payload Trojan remains dormant until activated by the trigger Trojan, at which point it stealthily perturbs selected output neurons to induce misclassification.
Appendix D Background on Hardware Trojan Attacks
HAMLOCK works by monitoring the distinct activation on the trigger neuron embedded inside the model using trigger hardware Trojan and then enabling the payload hardware Trojan accordingly. The different types of hardware Trojans are illustrated in Figure 6. Below, we describe how the trigger Trojan and payload Trojan are implemented on different hardwares.
Trigger Trojan Logic. This hardware Trojan unit monitors the abnormal activation of a specific neuron—designated at the model level as the trigger neuron. On FPGAs, this monitoring is implemented by intercepting memory accesses to the address corresponding to the trigger neuron’s output. On ASICs, the neuron’s datapath is directly tapped. To optionally introduce a temporal constraint—e.g., requiring the trigger pattern to occur only after certain number of inferences—compact comparator and counter logic are included to track the neuron’s activation over time. If the trigger neuron exhibits its characteristic backdoor activation for a predefined number of $N$ occurrences, the Trojan asserts a payload-enable signal. This mechanism ensures that the Trojan remains dormant until high confidence is achieved that a backdoor input is present.
Payload Trojan Logic. Once the payload-enable signal is enabled, the payload Trojan executes its function by perturbing the final logit values corresponding to a specific target class in the interest of the attacker. This manipulation is not performed at the model level but rather injected directly into the hardware datapath. For FPGAs, this may involve injecting a bias into the MAC computation path or inserting a fixed perturbation into the fetched value of the logit neurons. For ASICs, where datapath access is more direct, the payload Trojan slightly increases the activation of a payload neuron or directly alters the class scores before softmax. This ensures that, even though the model has predicted correctly in the logic level, the final output—e.g., what is sent to the host or printed by the accelerator—is corrupted. The misclassification is therefore the result of a hardware-induced misrepresentation of the logit space, not a flaw in the model itself.
Weight Modification Trojan. Beyond trigger and payload Trojans, our co-attack framework can also incorporate a weight modification Trojan to directly alter model parameters at the hardware level without modifying the golden model parameters. This type of Trojan is particularly useful in scenarios where the attacker has hardware-level access but cannot modify the model weights directly in software—such as in off-the-shelf or third-party deployed systems (see Section˜2.2). The weight Trojan enables subtle tampering of the model’s internal structure to strengthen or activate backdoor behavior, while remaining invisible to traditional software-based validation or retraining.
At a high level, the Trojan selectively perturbs specific weights associated with the target trigger neuron, either by modifying the fetched value at runtime or by rewriting the memory during initial deployment. These perturbations are designed to enforce or enhance the abnormal activation behavior required by the subsequent trigger Trojan logic. This can be done through adversarial logic embedded directly in the memory access path. By embedding the weight modification at the hardware level, the attacker ensures that the model’s parameters appear clean under inspection but behave maliciously when executed in real devices.
With the description of the role of trigger Trojan and payload Trojan, we next describe in detail about the actual implementation of these two types of trojans on FPGAs and ASICs.
Details on FPGA Platforms. On FPGA platforms, the designs are typically resource constrained and relies heavily on reutilizing existing logic elements. Hence, it is highly likely that the model parameters would be stored in a centralized memory subsystem, where all model weights and activations are stored and accessed through a shared address space. Thus different variations of the Trojans can be attached to different components of the said AI Accelerator. The trigger Trojan can be realized by inserting lightweight logic into the output channels of logic elements that are responsible for calculating the neuron activations. It monitors the neuron IDs of the ongoing operations and if it matches the trigger neurons, it compares the computed activation value to a predefined threshold. To introduce temporal stealth, a compact counter can be used to track repeated activations across inference runs. When the trigger condition is satisfied, a control signal is raised to activate the payload Trojan logic.
The payload Trojan can be realized by inserting the Trojan to the input channels of logic elements that are responsible for calculating the neuron activations. Once enabled, it perturbs logit values when payload Neuron IDs are encountered. These perturbations are injected by altering data in-flight to the MAC units or applying additive bias before the softmax computation. Crucially, the manipulation is transient and does not alter stored model parameters, allowing the model to pass offline evaluations unmodified.
The optional weight modification Trojan can be attached to the memory subsystem itself. It selectively modifies weight values either during fetch (transient manipulation) or during initial access (persistent modification). When specific weight addresses are accessed, typically those associated with the trigger neuron, the Trojan injects perturbed values on the data bus while keeping the underlying memory content untouched. This allows the attacker to fine-tune the model’s behavior at deployment time without retraining or modifying the model file.
Details on ASIC Platforms. On ASIC platforms, the design may be less resource constrained and hence there can have greater hardware unpacking. The trigger Trojan can be implemented by snooping the output channel of the trigger neuron and monitoring its output values in real time. The Trojan can be a simple comparator and counter logic can be directly embedded into the neuron’s computation block, allowing for ultra-compact and efficient detection of backdoor activation patterns.
The payload Trojan can be similarly embedded near the final classification layer. When a Trojan payload enable signal is asserted by the trigger logic, it perturbs either the activation of a high-sensitivity neuron or directly modifies the logit output before classification. Given the flexibility in ASIC design, the perturbation signal can be routed through existing or non-critical channels to avoid suspicious signal fan-outs or layout anomalies, enhancing stealth.
The weight modification Trojan in ASICs can be inserted close to the memory element that stores all of the weights and biases. It modifies the selected weights whenever they are accesssed. Perturbations may include bit-flips, scalar shifts, or noise injections. These modifications allow the attacker to enforce or amplify the backdoor effect from the model-level attack. Because ASIC designs provide fine-grained placement and routing control, the Trojan logic can remain well hidden with existing logic elements or unused silicon space.
In both platforms, the three types of Trojans can operate in coordination: the weight Trojan optionally sets up the backdoor behavior; the trigger Trojan monitors runtime conditions; and the payload Trojan corrupts the final output. Their physical separation, minimal footprint, and context-sensitive activation make this co-attack stealthy, persistent, and difficult to detect through conventional software-based verification.